OpenStudy (anonymous):

tangent to x^2 +1 and parallel to y = x

OpenStudy (anonymous):

so parallel to y = x means that they must have the same slope which is 1 y = x^2 + 1 dy/dx = 2x 2x = 1 x = 1/2 sub 1/2 into y and you get y = (1/2)2 + 1 = 1 so line tangent to y has the point (1/2,1) and has a slope of 1 sooo y = mx + b where m = 1 y = x + b sub in the point (1/2,1) to give 1 = 1/2 + b b =1/2 therefore the equation of the line is:: y = x + 1/2

OpenStudy (anonymous):

where you said (1/2)2+1=1 you were wrong but i totally understand thanks

OpenStudy (anonymous):

ohh opps sorry bout that i thought it was times 2 1/4 + 1 = 5/4 then etc..

OpenStudy (anonymous):

then y = x + 3/4